There are a total of 126 dogs and cats at a shelter.For every 5 dogs there are 4 cats. How many dogs are in the shelter?
Answer by stanbon(75887)

There are a total of 126 dogs and cats at a shelter.For every 5 dogs there are 4 cats. How many dogs are in the shelter?
---------------------
5x + 4x = 126 where 5x is # of dogs.
------
9x = 126
---
x = 14
-----
# of dogs = 5x = 5*14 = 70
--------
Cheers,
Stan H.
--------------------